OpenStudy (anonymous):

What 1/7 with binary. Explian.tku.

OpenStudy (amistre64):

binary fractions eh .... id have to read up on it

OpenStudy (anonymous):

that's a new one

OpenStudy (anonymous):

pretty sure i'm gonna pass on that

OpenStudy (anonymous):

there is an algorithm for this which i just looked up a week ago. but the idea is this. base two fractions look like \[\sum \frac{a_k}{2^k}\] where \[a_k=0,1\]

OpenStudy (anonymous):

so for example for 1/7, the largest "binary fraction" less than 1/7 is 1/8, so we are going to start with .001 meaning no 1/2,no 1/4, 1 1/8

OpenStudy (anonymous):

1/7=0,1428571429 in decimal

OpenStudy (anonymous):

then subtract and repeat. \[\frac{1}{7}-\frac{1}{8}=\frac{1}{56}\] so then next one will be .00101 meaning no 1/16 and 1 1/64 lather, rinse, repeat

OpenStudy (amistre64):

the website I reviewed says to take the decimal and multiply it by 2 and take the integer value of it: 1/7 = 7|1 = .1428 ...

OpenStudy (anonymous):

@amistre share the site if you would.

OpenStudy (anonymous):

0,1428571429*2=0,2857142857 0,2857142857*2=0,5714285714 0,5714285714*2=1,1428571429 0,1428571429*2=0,2857142857 0,2857142857*2=0,5714285714 0,5714285714*2=1,1428571429 ... so in binary 1/7 looks like = 0.001001001001001001...

OpenStudy (anonymous):

thanks makes it look much easier than the site a saw!

OpenStudy (anonymous):

wonder why my beginning was wrong though... \[\frac{1}{7}-\frac{1}{8}=\frac{1}{56}\] so why don't we have anything in the \[\frac{1}{64}\] place???

OpenStudy (anonymous):

oh because i am an idiot that's why

OpenStudy (anonymous):

1/64 = .000001

OpenStudy (amistre64):

i tried coping it and on a different terminal that went haywire :)

OpenStudy (anonymous):

let's check: \[\sum\frac{1}{8^k}=\frac{1}{1-\frac{1}{8}}=\frac{8}{7}\] subtract 1, get \[\frac{1}{7}\] fine

OpenStudy (amistre64):

pretty much it said; take the decimal form, multiply it by 2, and shave off the integer; move the decimal over and repeat

OpenStudy (anonymous):

right. and the answer is \[.001001001001...=\sum_{k=1}^{\infty}\frac{1}{8^k}\] which checks

OpenStudy (amistre64):

.0625 for instance would amount to: 2(.0625) = 0.1250 = 0 2(.625) = 1.25 = 1 2(.25) = 0.5 = 0 2(.5) = 1.0 = 1 .0101 perhaps?

OpenStudy (anonymous):

no you made a mistake i think

OpenStudy (amistre64):

its quite possible I did :)

OpenStudy (anonymous):

well first of all .0625 =1/16 right?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so it should be .0001

OpenStudy (anonymous):

second step you multiplied by 10!

OpenStudy (amistre64):

just going off my interp of the website :) ill have to review it again

OpenStudy (anonymous):

no i maean in your computation above look at step 2

OpenStudy (amistre64):

i misread the instructions :) i thought you moved the decimal over each time, but you actually just chop of the integer and work the new stuff 2(.0625) = 0.1250 = 0 2(.125) = 0.250 = 0 2(.25) = 0.5 = 0 2(.5) = 1.0 = 1 .0001

OpenStudy (anonymous):

that is it! but since 16 is a power of 2 that one was easy

OpenStudy (amistre64):

i like easy for testing ;)

OpenStudy (anonymous):

look what you did. you multiplied by 16, got 1 and said "oh it is 1/16=.0001

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

2(1/7) = 2/7 < 1 = 0 2(2/7) = 4/7 < 1 = 0 2(4/7) = 8/7 > 1 = 1 ; 1' 1/7 2(1/7) = 2/7 < 1 = 0 ....... .001001001...